How many sixteenths are there in 87 1/2%?
A)7
B)14
C)3.5
D)13
E)15

Answer by Alan3354(69443)

How many sixteenths are there in 87 1/2%?
A)7
B)14
C)3.5
D)13
E)15
----------------
87.5% = 87.5/100 = 7/8
7/8 = 14/16